# Flexing RISC-V Instruction Subset Processors (RISPs) to Extreme Edge

### v2

@@ 4.2 Power, Area and Frequency Analysis
We also introduce two baselines to compare the RISPs. The first baseline is an application-independent RISC-V processor supporting the full RV32E ISA (i.e.,  40 instructions) generated using the RISP methodology. This is called RISP-RV32E. The second baseline is Serv [39], - the world’s smallest 32-bit RISC-V processor. Serv is a bit serial processor supporting RV32I ISA, and its register file is mapped to on-chip memory rather than having a dedicated register file. Serv is configured to support RV32E (i.e., 16 general purpose registers rather than 32)
Figure 7: Energy per Instruction (EPI) in nanoJoules for RISPs and the two baselines (RISP-RV32E, and Serv)
@@
Figure 6(a) illustrates the maximum frequency (kHz) of RISPs compared to the two baselines. To determine the maximum operating frequency, we vary the input clock frequency while adjusting design constraints in the EDA tools. Starting at 100 kHz with significant positive slack, the frequency was incremented by 25 kHz steps until reaching 3 MHz, where the design became over-constrained with negative slack. The highest frequency with positive slack is identified as the maximum, and its corresponding area equivalence and power consumption are recorded for analysis. The maximum clock frequencies of RISPs range from 1,625 to 1,925 kHz, whilst RISP-RV32E and Serv can achieve up to 1,800 kHz and 1,950 kHz, respectively.
Figure 6(b) and Figure 6(c) present the NAND2-equivalent gatecount and power of the RISP for each application and compare them against RISP-RV32E and Serv, respectively. Across the complete set of the target applications, the NAND2-equivalent gatecount of the RISPs are 1,865-to-3,652. RISP-RV32E and Serv have 3,870 and 1,896, respectively. The power consumption of the RISPs has a range of 0.6-to-1.2 mW while RISP-RV32E and Serv consume 1.3 mW. Although Serv has a lower gate count compared to RISP-RV32E, its power consumption is nearly the same. This is attributed to the significantly higher proportion of flip-flops in Serv. In Serv, flip-flops account for approximately 52% of the total gates, whereas in RISP-RV32E, they constitute only 5% of the total area. The higher flip-flop density in Serv leads to increased power consumption.
4.3 Performance and Energy Analysis
As discussed in Section 3.3, RISPs consist of combinational elements except for the register that stores the Program Counter. Each instruction in the single-cycle RISP is executed in one clock cycle resulting in a Clock Cycles per Instruction (CPI) of one. The cycle time is set by the critical path of the RISP. An analysis of the achieved frequency of RISPs has been presented in Figure 6(a). It is important to note that this work assumes that the instruction fetch, data accesses and register file accesses are completed in the same clock cycle that are initiated.
In terms of energy efficiency, Figure 7 illustrates the Energy per Instruction (EPI) for the generated RISPs, RISP-RV32E and Serv. The x-axis represents the generated RISPs corresponding to each target application. For the calculation of the EPI, we assume a CPI of one for RISPs and RISP-RV32E. Serv is a bit-serial processor and has a CPI of thirty-two. The frequency and power values were obtained from Figures 6(a) and 6(c) for each processor, respectively. The results in Figure 7 highlight that RISPs can achieve at least 29 times better energy efficiency per instruction compared to Serv. The bit-serial architecture of Serv leads to a high CPI that substantially increases the number of cycles required per instruction, which results in much higher energy consumption per instruction.
4.4 Comparison of Extreme Edge RISPs Implemented as FlexICs
Figure 8: Layouts of the five implemented processors are shown. The leftmost layout corresponds to RISP-RV32E, used as one of the baselines. Next are the three RISPs for af_detect, armpit, and XGBoost — each representing an extreme-edge application. The rightmost layout is Serv, the second baseline. For each layout, the X and Y dimensions of the die area are shown in $\mu$m, along with the die area (in mm²) on the bottom-left of each layout. The total power consumption (in mW) is indicated on the bottom-right. Additionally, for the three RISPs, the number of independent instructions implemented in each design is annotated.
@@ 4.4 Comparison of Extreme Edge RISPs Implemented as FlexICs
This section presents full physical implementation of the three extreme edge RISPs as FlexICs and compare them to full physical implementation of the two baselines. RISPs are generated for the instruction subsets from the applications that are compiled with the --O0 flag. The physical implementation goes beyond the hardware synthesis of the three extreme edge RISPs and the baselines presented in Section 4.2 to develop the full-fledged layouts by floor planning, clock tree and buffer insertion, and place and route. The main goal for this is to show that the full physical implementation-related overheads (e.g., inserted buffers) do not hide the advantages of the RISP philosophy (i.e., not supporting unnecessary instructions in the processor is observably beneficial).
@@
The physical implementation is initially started at the maximum clock frequency of each processor found at the synthesis stage. Due to clock tree insertion and routing delays, the implementation cannot achieve the maximum clock frequency for all processors. We follow an iterative implementation process to incrementally reduce the clock frequencies until all processors became functional at 300 kHz. The implementation of all processors is done at this frequency using typical and nominal corner conditions with a supply voltage of 3V.
As shown in Figure 8, the die area utilization is significantly reduced in RISPs. The two baselines RISP-RV32E and Serv occupy 1.15 mm² and 0.83 mm², respectively. af_detect RISP has a die area of 0.83 mm², which is the same as Serv, yet it achieves a  28% reduction compared to RISP-RV32E. armpit and XGBoost RISPs achieve even further reductions, each occupying only 0.66 mm² — representing  21% savings over Serv and  43% over RISP-RV32E.
The two baselines RISP-RV32E and Serv have a total power consumption of 0.98 mW and 1.07 mW, respectively. In contrast, af_detect, armpit and XGBoost RISPs consume 0.91 mW, 0.69 mW and 0.81 mW, respectively, indicating noticeable power improvements in RISPs.
@@ 5 Generative AI based Code Retargeting
Figure 9: Generative AI based tool to implement code retargeting for RISPs used in long-lasting extreme edge applications
@@
Although the concept behind the RISP generation is to rely on the RISC-V software eco-system rather than generating a custom software toolchain, long-lasting extreme edge applications need occasional software updates because the applications may change. In this case, the applications need to be recompiled with possibly different compiler optimization flags, which may generate new instructions that are not supported by the RISP under consideration.
@@
There are two principal directions in which this question can be addressed. The first focuses on the code generation of the compiler backend that takes the instruction subset implemented in the RISP and constrains the code generation to this subset only. This method requires the modification of instruction selection and code generation phases of the compiler backend, which is not trivial. The second direction relies on the assumption that the code is generated for the full ISA and a separate tool is developed to take the full ISA code and re-compile it to the instruction subset supported by the RISP.
@@
We follow the second approach where we develop a tool that transforms the code generated by the compiler targeting the full RISC-V ISA (e.g., riscv32-unknown-elf-gcc) to the code that use only the instructions in the target subset using Generative AI. This is a significant step towards re-usability of RISPs after fabrication. Figure 9 presents the retargeting framework along with a simple example. Given an application and a target-subISA, the tool is able to identify the instructions that are not included in the target-subISA. Then, with the use of ChatGPT RISC-V Assembly plugin [8], every instruction that is not part of the target-subISA is transformed to a set of instructions that are included in the subISA and is defined as an independent macro. The macros of the transformed instructions are stored in a macro.S file. Finally, the application is recompiled using the standard riscv32-unknown-elf-gcc with the macro.S file.
@@
As mentioned before, for the transformation of the instructions that are not part of the subISA, we use the ChatGPT RISC-V Assembly plugin. Initially, we provide the target-subISA and the instructions that need to be transformed. Objectives that restrict the solution space can be added based on the user preferences. For example, such objectives could be “allow the use of temporary registers to provide a solution”, or, “keep the codesize to minimum as you transform the instructions”. After that, the generation of macros for the instructions begins and the LLM provides a solution. Then, the macros are verified against the original semantics of the transformed instruction with custom test cases. In the case of failure, we notify the LLM and the process is initiated again, until we reach a functionally verified set of macros.
@@
To evaluate the above framework, we define a minimal instruction subset of twelve instructions {addi, add, and, xori, sll, sra, jal, jalr, blt, bltu, lw, sw} from which other instructions can reproduced. armpit can be defined as a “long-lasting extreme edge application” and for this experiment it is chosen as a case study to show the potential of the framework. The framework is provided with the vanilla version of the target application compiled with -O2, and we instruct the LLM to transform the instructions of the armpit application to the minimal instruction subset. Our results show that the code size after the re-compilation of the application is 11.6% larger than the original version of the application. This experiment demonstrates the feasibility of this approach to efficiently re-target code but with an overhead of larger code size and slower runtime of the application.
@@ 6 Conclusions
In this paper, we have proposed a methodology for generating automatically a 32-bit RISC-V instruction subset single-cycle processor, called RISP, implemented as FlexICs. The methodology constructs a RISP supporting only the instruction set identified during the characterization of an extreme edge application or domain. Initially, the semantics of each instruction in the RISC-V ISA are translated into a formally verified and discrete, hardware block and stored in a library. The hardware blocks of the identified instruction set are extracted from the library, stitched together to build the core of the RISP, which is then connected to the rest of the pre-verified processor components to generate the RISP. The RISP generation methodology reduces both design and verification time by making the verification a first-class constraint in the design time. Our experimental results show that RISPs are, on average, 30% more power- and area-efficient than a RISC-V processor supporting the full instruction set for a range of extreme edge and embedded applications, and are nearly 30 times more energy efficient with respect to Serv - the world’s smallest 32-bit RISC-V processor
The proposed methodology can be extended to generate pipelined RISPs if higher clock frequencies are required. However, it is not suitable for generating complex processors that have deeper superscalar and out-of-order execution pipelines. Extreme edge applications do not require such high performance processors, so our focus has been to develop a microcontroller-class RISP that is a single-cycle scalar in-order execution machine providing adequate performance for applications at extreme edge.

### v3

@@ 4.2 FlexIC Synthesis Results
The RTL of each RISSP is synthesized using Pragmatic’s 0.6µm IGZO-based FlexIC process using a commercial EDA tool. Each RISSP is synthesized without the RF to better understand the effects of the instruction subsets in the hardware. We also introduce two baselines to compare the RISSPs. The first baseline is an application-independent RISC-V processor supporting the full RV32E ISA (~40 instructions) generated using the RISSP methodology. This is called RISSP-RV32E. The second baseline is Serv [38], - the world’s smallest 32-bit RISC-V processor. Serv is a bit serial processor supporting RV32I ISA, and its RF is mapped to on-chip memory rather than having a dedicated RF. Serv is configured to support RV32E (i.e., 16 general purpose registers rather than 32).
Figure 6: Maximum clock frequencies at which RISSPs, RISSP-RV32E and Serv that can operate
4.2.1 Maximum Frequency Analysis
@@
Figure 6 illustrates the maximum frequency (kHz) of RISSPs compared to the two baselines. To determine the maximum operating frequency, we vary the input clock frequency while adjusting design constraints in the EDA tools. Starting at 100 kHz with significant positive slack, the frequency was incremented by 25 kHz steps until reaching 3 MHz, where the design became over-constrained with negative slack. The highest frequency with positive slack is identified as the maximum. The maximum clock frequencies of RISSPs range from 1,500 to 1,850 kHz, whilst RISSP-RV32E and Serv can achieve up to 1,700 kHz and 2,050 kHz, respectively.
4.2.2 Area Analysis
Figure 7 presents the average area of the RISSP for each application and the baselines. The average area is calculated from the NAND2-equivalent gatecounts of the processor across the range of frequencies with positive slack. RISSPs occupy 8-to-43% less area compared to RISSP-RV32E. The smallest RISSP (i.e., xgboost) is 23% larger than Serv.
Figure 7: Average NAND2-equivalent gate count of RISSPs compared to Serv and RISSP-RV32E across the range of frequencies with positive slack
Figure 8: Average power (static + dynamic) comparison of the RISSPs, Serv and RISSP-RV32E across the range of frequencies with positive slack
4.2.3 Power Analysis
Figure 8 shows the average power of the RISSP for each application and the baselines. The average powered is calculated in a similar manner to Section 4.2.2. RISSPs consume 3-to-30% less power than RISSP-RV32E. Although Serv has a lower gatecount compared to RISSPs and RISSP-RV32E, its power consumption is higher than them. For example, it consumes 40% more power than RISSP-RV32E. This is mainly because Serv has a higher proportion of flip-flops (FFs) compared to the RISSP microarchitecture, and an FF in the FlexIC process consumes 10 times more power than a NAND2 gate.
4.2.4 Energy Efficiency
Figure 9 illustrates the Energy per Instruction (EPI) for the RISSPs, RISSP-RV32E and Serv. The x-axis represents the RISSPs corresponding to each target application. Because a RISSP is a single-cycle processor, the clock cycles per instruction (CPI) for RISSPs and RISSP-RV32E is 1. On the other hand, Serv is a bit-serial processor, and has a CPI of 32 on average. The EPI is calculated by dividing the power consumption of the design point at the maximum frequency by that frequency and multiply the product by the CPI. The results highlight that RISSP-RV32E and RISSPs are about 35 and 40 times more energy-efficient than Serv. The bit-serial architecture of Serv leads to a high CPI resulting in much higher energy consumption per instruction.
Figure 9: Energy per Instruction (EPI) in nanoJoules for RISSPs and the two baselines (RISSP-RV32E, and Serv)
4.3 FlexIC Physical Implementation Results
@@
We present full physical implementation of the three extreme edge RISSPs as FlexICs and compare them to full physical implementation of the two baselines. RISSPs are generated for the instruction subsets from the applications compiled with the --O2 flag. The physical implementation goes beyond hardware synthesis to full-fledged FlexIC layouts by floorplanning, clock tree insertion, and place and route. The goal is to show that the full physical implementation-related overheads (e.g., inserted buffers) do not hide the advantages of the RISSP philosophy (i.e., not supporting unnecessary instructions in the processor is observably beneficial).
Figure 10: Layouts of the five implemented processors are shown. The leftmost layout corresponds to RISSP-RV32E. Next are the three RISSPs for af_detect, armpit, and xgboost — each representing an extreme-edge application. The rightmost layout is Serv. For each layout, the X and Y dimensions of the die area are shown in µm, along with the die area (in mm²) and percentage of FFs being used in each design, on the bottom-left of each layout. The total power consumption (in mW) is indicated on the bottom-right. Additionally, for the three RISSPs, the number of distinct instructions implemented in each design is annotated.
@@
The physical implementation is initially started at the maximum clock frequency of each processor found at the synthesis stage. Due to clock tree insertion and routing delays, the implementation cannot achieve the maximum clock frequency for all processors. We follow an iterative implementation process to incrementally reduce the clock frequencies until all processors became functional at 300 kHz. All processors are implemented at this frequency using typical and nominal corner conditions with a supply voltage of 3V.
Figure 10 shows the layouts of the three RISSPs, RISSP-RV32E and Serv. RISSP-af_detect achieves an  8% reduction in area compared to RISSP-RV32E whilst RISSP-armpit has the same area as Serv but reduces the area by ~35% with respect to RISSP-RV32E. RISSP-xgboost achieves even further reduction by ~11% and ~42% over Serv and RISSP-RV32E. This is because Serv has a high proportion of FFs (i.e., 60%), and therefore requires more clock tree buffers, which increases its area despite the fact that the synthesis results indicating that Serv is smaller than RISSP-xgboost.
The two baselines RISSP-RV32E and Serv have almost the same power consumption in spite of Serv being ~35% smaller than RISSP-RV32E. This is again because of the high proportion of FFs in Serv whereas FFs take up only ~6% of RISSP-RV32E. RISSP-af_detect consume almost the same power as Serv and RISSP-RV32E whilst RISSP-armpit and RISSP-xgboost consume ~8% and ~21% less power than Serv and RISSP-RV32E, respectively.
Figure 11: Generative AI based tool to implement code retargeting for RISSPs used in long-lasting extreme edge applications
@@ 5 Generative AI based Code Retargeting
Although the concept behind the RISSP generation is to rely on the RISC-V software ecosystem rather than generating a custom software toolchain, long-lasting extreme edge applications need frequent software updates because the target applications may change. For instance, applications need to be recompiled with different compiler optimization flags, which may generate new instructions that are not supported by the RISSP under consideration.
@@
There are two possible approaches to address this scenario. The first approach focuses on modifications at the compiler backend that needs to be modified to receive the instruction subset implemented in the RISSP and constrain the assembly code generation to this subset only. This method is not trivial and requires significant changes in the instruction selection and code generation phases of the compiler backend. The second approach relies on the assumption that the assembly code is generated using standard compiler toolflows and a separate tool is responsible to transform the generated assembly into a version that uses the instruction subset supported by the RISSP.
@@
We follow the second approach where we develop a tool that uses generative AI to transform the assembly code generated by the compiler targeting the full RISC-V ISA (e.g., riscv32-unknown-elf-gcc) to assembly code that uses only the instructions in the instruction subset. This is an important step towards reusability of RISSPs after fabrication. Figure 11 presents the retargeting framework along with a simple example. Given an application and an instruction subset, the tool is able to identify the instructions that are not included in the instruction subset. Then, with the use of ChatGPT RISC-V Assembly plugin [8], every instruction that is not part of the instruction subset is transformed into a set of instructions that are included in the instruction subset and is defined as an independent macro. The macros of the transformed instructions are stored in a macro.S file. Finally, the application is recompiled using the standard riscv32-unknown-elf-gcc with the macro.S file.
@@
Initially, we provide the instruction subset and the instructions that need to be transformed to the ChatGPT RISC-V Assembly plugin. Objectives that restrict the solution space can be added based on the user preferences. For example, such objectives could be “allow the use of temporary registers to provide a solution”, or, “keep the codesize to minimum as you transform the instructions”. After that, the process of generating macros for the non-supported instructions from the instruction subset begins. Then, the macros are verified against the original semantics of the transformed instruction with custom test cases. If the LLM generates a macro that cannot be functionally verified, the macro is rejected, and then another macro is requested from the LLM. This continues until a valid macro is generated. In our experiments, a valid macro can be generated on less than 10 attempts.
Figure 12: Comparison of code size (bytes) and the number of distinct instructions between the initial binary and the retargeted binary using LLM
@@
To evaluate the above framework, we define a minimal instruction subset of twelve instructions {addi, add, and, xori, sll, sra, jal, jalr, blt, bltu, lw, sw} from which other instructions can be reproduced. Armpit, xgboost and af_detect can be defined as “long-lasting extreme edge applications”, as discussed in Section 4, and for this experiment they are chosen as a case study to show the potential of the framework. The framework receives as input the initial version of the target applications compiled with -O2, and attempts to transform the non-supported instructions of the target applications to functionally relevant macro based on the instruction subset.
Figure 12 shows that the code size after the recompilation of the applications are 13%, 5.2% and 36% larger than the original version for armpit, xgboost and af_detect, respectively. The number of distinct instructions shows minimal variation for armpit and xgboost, while the number of distinct instructions for af_detect is reduced from 23 to 12. This experiment demonstrates the feasibility of this approach to efficiently retarget code but with an overhead of larger code size and potentially slower runtime of the application.
@@ 6 Conclusions
We have proposed a methodology for automatically generating a single-cycle 32-bit RISC-V instruction subset processor (RISSP) to be deployed as a FlexIC to meet the computing requirements of the extreme edge. The methodology constructs a RISSP supporting only the instruction subset identified during the characterization of an extreme edge application or domain. The semantics of each instruction in the RISC-V ISA are translated into a formally verified and discrete, hardware block and stored in a library. The hardware blocks of the instruction subset are extracted from the library, stitched together to build the core of the RISSP, which is then connected to the rest of the pre-verified processor components to generate the RISSP. The RISSP methodology reduces the time-to-market of custom processors by integrating verification as a primary design principle. The synthesis results have shown that RISSPs can achieve 8-to-43% reduction in area and 3-to-30% reduction in power compared to a processor supporting the full RISC-V ISA, and are on average about 40 times more energy efficient than Serv- the world’s smallest 32-bit RISC-V processor. When implemented as FlexICs, the RISSPs targeting extreme edge applications can achieve up to 42% and 21% less area and power.
The proposed methodology can support custom instructions that can be added to the pre-verified ISA hardware library like other instructions in the ISA. Also, the methodology can be extended to generate pipelined RISSPs if higher clock frequencies are required but is not suitable for generating complex processors that have deeper superscalar and out-of-order execution pipelines. Extreme edge applications do not require such high performance processors, so our focus has been to develop a microcontroller-class RISSP that is a single-cycle scalar in-order execution machine providing adequate performance for applications at extreme edge.
7 Acknowledgments
This work is partially funded by GRASP EU Project No 101134936. The authors would like to thank John Biggs, a co-founder of Arm and a Pragmatic fellow, for his valuable inputs on physical implementation analysis, and also Bruce Hoult from the RISCV sub-reddit forum for his comments on macro retargeting.